# SARDP_supp

### v1

@@ 3. THEORETICAL PROPERTIES
In this section, we theoretically analyze both the numerical and statistical properties of the CLE and CLS. First, we introduce some theoretical assumptions.
@@ 3.1. Theoretical Assumptions
Define $\|B\|_{1}=\max_{j}\sum_{i}|b_{ij}|$ to be $L_{1}$ norm and $\|B\|_{\infty}=\max_{i}\sum_{j}|b_{ij}|$ to be $L_{\infty}$ norm for an arbitrary matrix $B=(b_{ij})\in\mathbb{R}^{N\times N}$. For theoretical analysis, the following regularity conditions are required.
@@
(C3) (Noise Term) Assume that all the ${e_{i}}$s, $\epsilon_{i}$s, and $\epsilon_{ij}^{x}$s ($1\leq i,j\leq N$) are independently and identically distributed with mean zero. For the variance, we assume that $\text{Var}(\epsilon_{i})=\lambda^{2}$ and $\text{Var}(\epsilon_{ij}^{x})=\lambda_{x}^{2}$. To simplify the asymptotic covariance form of the estimators, we assume $E(e_{i}^{3})=E(\epsilon_{i}^{3})=E(\epsilon_{ij}^{x3})=0$. For higher moment conditions, we assume that $E(e_{i}^{4})=\mu_{4}^{e}$, $E(\epsilon_{i}^{4})=\mu_{4}^{\epsilon}$, $E(\epsilon_{ij}^{x4})=\mu_{4}^{\epsilon_{x}}$, $E(e_{i}^{2}-\sigma_{0}^{2})^{4}=c^{\prime}_{e}$, $E(\epsilon_{i}^{2}-\lambda^{2})^{4}=c^{\prime}_{\epsilon}$, and $E(\epsilon_{ij}^{x2}-\lambda_{x}^{2})^{4}=c^{\prime}_{\epsilon_{x}}$, for positive constants $\mu_{4}^{e}$, $\mu_{4}^{\epsilon}$, $\mu_{4}^{\epsilon_{x}}$, $c^{\prime}_{e},$ $c^{\prime}_{\epsilon}$, and $c^{\prime}_{\epsilon_{x}}$.
Condition (C1) requires the covariance to exist and be non-singular. The same condition was adopted by Lee (2004), Yang and Lee (2017), and Zhu et al. (2020). Condition (C2) requires the uniform boundedness of features $\mathbb{X}$, the weighting matrix $W$, and $S_{0}^{-1}$; this is a classical regularity condition in the SAR model (Lee, 2004; Yang and Lee, 2017). Condition (C3) is a typical assumption for the noise term (Huang et al., 2019; Zhu et al., 2020; Huang et al., 2020). It is remarkable that the artificially created noise is allowed to be non-Gaussian, as long as the moment conditions are satisfied. Furthermore, condition $E(e_{i}^{3})=E(\epsilon_{i}^{3})=E(\epsilon_{ij}^{x3})=0$ is assumed for convenience. The aim is to simplify the form of the asymptotic covariances for the estimators. Notably, the theory to be presented can be softly generalized under the assumption of non-zero third-order moments using the same theoretical proof techniques in this work. Given these conditions, we then establish the theoretical results in the subsequent subsections.
@@ 3.2. Theoretical Properties of the CLE
However, to estimate the parameters for the PSAR model based on the likelihood function, we need to analyze $\Omega^{-1}=\big{(}\sigma^{2}I_{N}+\lambda^{2}SS^{\top}\big{)}^{-1}$, which distinguishes the estimation procedure from the previous two approaches. On the one hand, even if $W$ and $S^{-1}$ satisfy the boundedness conditions of the $L_{1}$ and $L_{\infty}$ norms, it cannot be directly proven that $\Omega^{-1}$ is bounded in $L_{1}$ and $L_{\infty}$ norms. Consequently, the first method cannot be employed. On the other hand, notice that $S^{\top}S$ in $\Omega^{-1}$ cannot be analyzed as a transition-probability matrix; thus, the assumption of the existence of a stationary distribution cannot be made either. This makes the second method inapplicable. Therefore, we need to reexamine the theoretical properties of the CLE $\hat{\theta}_{\text{CLE}}$ based on the PSAR model. We begin with the discussion of the identification issue.
The aim of the discussion for identification is to demonstrate that $\theta_{0}$ could be identified by equations (2.6)–(2.8). We then verify that this is equivalent to clarifying that $\theta_{0}$ can be identified by minimizing the negative log-likelihood function $E\big{\{}\mathcal{L}(\theta)\big{\}}$ defined in (2.4) for the observed response $\mathbb{Y}^{*}$. To this end, we define
${\bf Q}_{\rm CL}(\theta)=\mathcal{L}^{*}(\theta)-\frac{1}{2}\lambda^{2}_{x}\beta_{2}\beta_{2}^{\top}\mbox{tr}\big{(}\Omega^{-1}\big{)}$
to be one of the primitive functions of ˙$\dot{\mathcal{L}}^{*}(\theta)-\Delta\mathbf{S}_{\text{CL}}(\theta)$, which means that $\dot{\textbf{Q}}_{\text{CL}}(\theta)=\dot{\mathcal{L}}^{*}(\theta)-\Delta\mathbf{S}_{\text{CL}}(\theta)$. Subsequently, define $\mathcal{Q}_{\text{CL}}(\theta)=E\{{\bf Q}_{\text{CL}}(\theta)\}$. Then, we have,
$\mathcal{Q}_{\rm CL}(\theta)=-\log|S|+\frac{1}{2}\log|\Omega|+\frac{1}{2}\mbox{tr}\big{\{}S^{\top}\Omega^{-1}SS_{0}^{-1}\Omega_{0}(S_{0}^{-1})^{\top}\big{\}}+\frac{1}{2}\beta^{\top}\mathbb{X}^{\top}\Omega^{-1}\mathbb{X}\beta$
$~{}~{}~{}~{}+\frac{1}{2}\beta_{0}^{\top}\mathbb{X}^{\top}(S_{0}^{-1})^{\top}S^{\top}\Omega^{-1}SS_{0}^{-1}\mathbb{X}\beta_{0}-\beta^{\top}\mathbb{X}^{\top}\Omega^{-1}SS_{0}^{-1}\mathbb{X}\beta_{0}=E\big{\{}\mathcal{L}(\theta)\big{\}},$
and $\dot{\mathcal{Q}}_{\rm CL}(\theta)=E\big{\{}\dot{\mathcal{L}}^{*}(\theta)\big{\}}-\Delta\mathbf{S}_{\rm CL}(\theta)$. This means that solving equations (2.6)–(2.8) is equivalent to minimizing $\mathcal{Q}_{\rm CL}(\theta)$. Consequently, to address the identification issue, it suffices to verify that $\theta_{0}$ can be identified by minimizing $\mathcal{Q}_{\rm CL}(\theta)=E\{\mathcal{L}(\theta)\}$.
Next, note that for any parameter $\theta$ in the parameter space, by the calculation of (2.7), we can express $\beta$ as a function of $\rho,\sigma^{2}$ and $\theta_{0}$ as,
$\beta(\rho,\sigma^{2})=(\mathbb{X}^{\top}\Omega^{-1}\mathbb{X})^{-1}\mathbb{X}^{\top}\Omega^{-1}SS^{-1}_{0}\mathbb{X}\beta_{0}.$ (3.1)
Consequently, one only needs to verify the identification of $\rho_{0}$ and $\sigma_{0}^{2}$, which will lead to the identification of $\theta_{0}$. By substituting (3.1) into $\mathcal{Q}_{\rm CL}(\theta)$, we have
$\mathcal{Q}_{\rm CL}(\rho,\sigma^{2})$ $=$ $-\log|S|+\frac{1}{2}\log|\Omega|+\frac{1}{2}\beta_{0}^{\top}\mathbb{X}^{\top}(S_{0}^{-1})^{\top}S^{\top}\Omega^{-1}H_{\Omega}SS_{0}^{-1}\mathbb{X}\beta_{0}$ (3.2)
$+\frac{1}{2}\mbox{tr}\big{\{}S^{\top}\Omega^{-1}SS_{0}^{-1}\Omega_{0}(S_{0}^{-1})^{\top}\big{\}},$
where $H_{\Omega}=I_{N}-\mathbb{X}(\mathbb{X}^{\top}\Omega^{-1}\mathbb{X})^{-1}\mathbb{X}^{\top}\Omega^{-1}$. Then, to address the identification issue, we only need to demonstrate that $(\rho_{0},\sigma_{0})$ can be identified by minimizing (3.2). Before presenting the theoretical results, the following identification condition is required.
@@
(Identification of CLE) Assume (C1)–(C4) hold. Then, $(\rho_{0},\sigma_{0}^{2})$ is identifiable. That is, for any $\epsilon>0$, we have
$\liminf_{N\to\infty}\min_{(\rho,\sigma^{2})\in\bar{N}_{\epsilon}(\rho_{0},\sigma_{0}^{2})}\frac{1}{N}\big{\{}\mathcal{Q}_{\rm CL}(\rho,\sigma^{2})-\mathcal{Q}_{\rm CL}(\rho_{0},\sigma_{0}^{2})\big{\}}>0.$
Here, $\bar{N}_{\epsilon}(\rho_{0},\sigma_{0}^{2})$ is the complement of an open neighborhood of $(\rho_{0},\sigma_{0}^{2})$ in parameter space and has a diameter of $\epsilon$.
The proof of Theorem 1 is given in Appendix C.1. The theorem guarantees the identification of $(\rho_{0},\sigma_{0}^{2})$, which subsequently leads to the identification of $\theta_{0}$. We then examine the statistical property of the CLE, which results in the following theorem.
@@ 3.3. Discussion of Statistical Inference for the CLE
Remarkably, for the inference of $\hat{\theta}_{\text{CLE}}$, the asymptotic covariance in (3.3) cannot be immediately obtained when $\mathbb{X}$ is unobserved. Therefore, we denote $\Sigma_{k}^{\text{CL}}=\Sigma_{k}^{\text{CL}}(\mathbb{X},\theta)$ $(k=1,2)$ as a function of $\mathbb{X}$ and $\theta$. In this subsection, for simplicity, we use $\theta$ to represent the true parameter. Then, the difference between $\Sigma_{k}^{\text{CL}}(\mathbb{X}^{*},\theta)$ and $\Sigma_{k}^{\text{CL}}(\mathbb{X},\theta)$ should be considered. This is because data users can only observe $\mathbb{X}^{*}$ instead of $\mathbb{X}$, and statistical inference can be made relying only on $\mathbb{X}^{*}$. Thus, define $\Delta_{k}(\theta)=E\{\Sigma_{k}^{\text{CL}}(\mathbb{X}^{*},\theta)\}-\Sigma_{k}^{\text{CL}}(\mathbb{X},\theta)$, and $G=WS^{-1}$. Then, we have
$\Delta_{1}^{\rm CL}(\theta)$ $=$ $\lambda_{x}^{4}\beta_{2}^{\top}\beta_{2}\begin{pmatrix}\beta_{2}^{\top}\beta_{2}\mbox{tr}(G^{\top}\Omega^{-2}G)&{\bf 0}_{p_{1}}^{\top}&\mbox{tr}(\Omega^{-2}G)\beta_{2}^{\top}&0\\ {\bf 0}_{p_{1}}&{\bf 0}_{p_{1}\times p_{1}}&{\bf 0}_{p_{1}\times p_{2}}&{\bf 0}_{p_{1}}\\ \mbox{tr}(\Omega^{-2}G)\beta_{2}&{\bf 0}_{p_{2}\times p_{1}}&\mbox{tr}(\Omega^{-2})I_{p_{2}}&{\bf 0}_{p_{2}}\\ 0&{\bf 0}_{p_{1}}^{\top}&{\bf 0}_{p_{2}}^{\top}&0\end{pmatrix},$
$\Delta_{2}^{\rm CL}(\theta)$ $=$ $\lambda_{x}^{2}\begin{pmatrix}\beta_{2}^{\top}\beta_{2}\mbox{tr}(G^{\top}\Omega^{-1}G)&{\bf 0}_{p_{1}}^{\top}&\mbox{tr}(\Omega^{-1}G)\beta_{2}^{\top}&\beta_{2}^{\top}\beta_{2}\mbox{tr}(\Omega^{-2}G)\\ {\bf 0}_{p_{1}}&{\bf 0}_{p_{1}\times p_{1}}&{\bf 0}_{p_{1}\times p_{2}}&{\bf 0}_{p_{1}}\\ \mbox{tr}(\Omega^{-1}G)\beta_{2}&{\bf 0}_{p_{2}\times p_{1}}&\mbox{tr}(\Omega^{-1})I_{p_{2}}&{\bf 0}_{p_{2}}\\ \beta_{2}^{\top}\beta_{2}\mbox{tr}(\Omega^{-2}G)&{\bf 0}_{p_{1}}^{\top}&{\bf 0}_{p_{2}}^{\top}&0\end{pmatrix}.$
Therefore, based on the idea of bias correction, we can obtain a bias-corrected estimator for each element of the covariance matrix in (3.3) based on the observed $\mathbb{X}^{*}$. The estimators can be denoted as $\widehat{\Sigma}_{k}^{\text{CL}}=\Sigma_{k}^{\text{CL}}(\mathbb{X}^{*},\hat{\theta}_{\text{CLE}})-\Delta_{k}^{\text{CL}}(\hat{\theta}_{\text{CLE}})$ $(k=1,2)$ with $\hat{\theta}_{\text{CLE}}$ plugged in. Employing a technique similar to that used in the proof of Lemma 6 in Appendix B, the consistency of $\widehat{\Sigma}_{k}^{\text{CL}}$ can be established. Then, the estimated standard error of $\hat{\theta}_{\text{CLE}}$ could be calculated by the estimated asymptotic covariance matrix $(\widehat{\Sigma}_{2}^{\text{CL}})^{-1}\widehat{\Sigma}_{1}^{\text{CL}}(\widehat{\Sigma}_{2}^{\text{CL}})^{-1}$.
@@ 3.4. Theoretical Properties of the CLS
Here, the identification issue is discussed in two cases: with covariates and without covariates. The condition is similar to that in Zhu et al. (2020) and Huang et al. (2020). We then have the following theorem.
@@
The proof of Theorem 4 is given in Appendix C.4. Next, we establish the following statistical theoretical properties of $\hat{\gamma}_{\text{CLS}}$.
@@
(Asymptotic Normality of CLS) Assume that (C1)–(C3) and (C4) hold. We then have $\sqrt{N}(\hat{\gamma}_{\rm CLS}-\gamma_{0})\to_{d}N(\mathbf{0}_{p+1},(\Sigma_{2}^{\rm LS})^{-1}\Sigma_{1}^{\rm LS}(\Sigma_{2}^{\rm LS})^{-1})$ as $N\to\infty$, where $\Sigma_{1}^{\rm LS}$ and $\Sigma_{2}^{\rm LS}$ are assumed to be positive definite matrices expressed as
@@
Through Theorem 5, we can conclude that the CLS estimator $\hat{\gamma}_{\text{CLS}}$, as well as the CLE discussed earlier, are both $\sqrt{N}$-consistent. Practically, we obtain $\hat{\gamma}_{\text{CLS}}^{(t)}$ by Algorithm 2 instead of $\hat{\gamma}_{\text{CLS}}$. Similar to the discussion of $\hat{\theta}_{\text{CLE}}^{(t)}$, we need to establish the numerical convergence of $\hat{\gamma}_{\text{CLS}}^{(t)}$. Therefore, the following theorem is presented.
Theorem 6.
(Numerical Convergence of CLS) Assume that (C1)–(C3) and (C4) hold. If the initial value $\hat{\gamma}_{\rm CLS}^{(0)}$ lies close to $\hat{\gamma}_{\rm CLS}$, then $\hat{\gamma}_{\rm CLS}^{(t)}\to\hat{\gamma}_{\rm CLS}$ as $t\to\infty$ with probability tending to 1.
For statistical inference, similar to the discussion for $\hat{\theta}_{\text{CLE}}$, we can also provide consistent estimators for each element in the asymptotic covariance matrix (3.4) based on the observed $\mathbb{X}^{*}$. Define $\Sigma_{k}^{\text{LS}}=\Sigma_{k}^{\text{LS}}(\mathbb{X},\gamma)$ $(k=1,2)$ as functions of $\mathbb{X}$ and $\gamma$, which are expressed in (3.4). Then, we could plug $\mathbb{X}^{*},\hat{\gamma}_{\text{CLS}}$ in to obtain $\Sigma_{k}^{\text{LS}}(\mathbb{X}^{*},\hat{\gamma}_{\text{CLS}})$ for $k=1,2$. As a result, $\Delta_{k}^{\text{LS}}(\gamma)=E\{\Sigma_{k}^{\text{LS}}(\mathbb{X}^{*},\theta)\}-\Sigma_{k}^{\text{LS}}(\mathbb{X},\theta)$ could be calculated as,
$\Delta_{1}^{\rm LS}(\gamma)$ $=$ $4\lambda_{x}^{4}\beta_{2}^{\top}\beta_{2}\begin{pmatrix}\beta_{2}^{\top}\beta_{2}\mbox{tr}\{G^{\top}(Sd_{\rho}^{2}S^{\top})^{2}G\}&{\bf 0}_{p_{1}}^{\top}&\mbox{tr}\{(Sd_{\rho}^{2}S^{\top})^{2}G\}\beta_{2}^{\top}\\ {\bf 0}_{p_{1}}&{\bf 0}_{p_{1}\times p_{1}}&{\bf 0}_{p_{1}\times p_{2}}\\ \mbox{tr}\{(Sd_{\rho}^{2}S^{\top})^{2}G\}\beta_{2}&{\bf 0}_{p_{2}\times p_{1}}&\mbox{tr}\{(S_{0}d_{\rho_{0}}^{2}S^{\top})^{2}\}I_{p_{2}}\end{pmatrix}$
$\Delta_{2}^{\rm LS}(\gamma)$ $=$ $2\lambda_{x}^{2}\begin{pmatrix}\beta_{2}^{\top}\beta_{2}\mbox{tr}(G^{\top}Sd_{\rho}^{2}S^{\top}G)&{\bf 0}_{p_{1}}^{\top}&\mbox{tr}(Sd_{\rho}^{2}S^{\top}G)\beta_{2}^{\top}\\ {\bf 0}_{p_{1}}&{\bf 0}_{p_{1}\times p_{1}}&{\bf 0}_{p_{1}\times p_{2}}\\ \mbox{tr}(Sd_{\rho}^{2}S^{\top}G)\beta_{2}&{\bf 0}_{p_{2}\times p_{1}}&\mbox{tr}(Sd_{\rho}^{2}S^{\top})I_{p_{2}}\end{pmatrix}.$

### v2

@@ 3. THEORETICAL PROPERTIES
In this section, we first theoretically analyze both the numerical and statistical properties of the CLE and CLS. Then, the generalizations of PSAR are carefully presented to extend the method to more model forms, enabling support for a wider range of data analysis applications. First, we introduce some theoretical assumptions.
@@ 3.1. Theoretical Assumptions
Define $\|B\|_{1}=\max_{j}\sum_{i}|b_{ij}|$ to be $L_{1}$ norm and $\|B\|_{\infty}=\max_{i}\sum_{j}|b_{ij}|$ to be $L_{\infty}$ norm for an arbitrary matrix $B=(b_{ij})$. For theoretical analysis, the following regularity conditions are required.
@@
(C3) (Noise Term) Assume that all the ${e_{i}}$s, $\varepsilon_{i}$s, and $\varepsilon_{ij}^{x}$s ($1\leq i,j\leq N$) are independently and identically distributed with mean zero. For the variance, we assume that $\text{Var}(\varepsilon_{i})=\lambda^{2}$ and $\text{Var}(\varepsilon_{ij}^{x})=\lambda_{x}^{2}$. To simplify the asymptotic covariance form of the estimators, we assume $E(e_{i}^{3})=E(\varepsilon_{i}^{3})=E(\varepsilon_{ij}^{x3})=0$. For higher moment conditions, we assume that $E(e_{i}^{4})=\mu_{4}^{e}$, $E(\varepsilon_{i}^{4})=\mu_{4}^{\varepsilon}$, $E(\varepsilon_{ij}^{x4})=\mu_{4}^{\varepsilon_{x}}$, $E(e_{i}^{2}-\sigma_{0}^{2})^{4}=c^{\prime}_{e}$, $E(\varepsilon_{i}^{2}-\lambda^{2})^{4}=c^{\prime}_{\varepsilon}$, and $E(\varepsilon_{ij}^{x2}-\lambda_{x}^{2})^{4}=c^{\prime}_{\varepsilon_{x}}$, for positive constants $\mu_{4}^{e}$, $\mu_{4}^{\varepsilon}$, $\mu_{4}^{\varepsilon_{x}}$, $c^{\prime}_{e},$ $c^{\prime}_{\varepsilon}$, and $c^{\prime}_{\varepsilon_{x}}$.
Condition (C1) requires the covariance to exist and be non-singular. The same condition was adopted by Lee (2004), Yang and Lee (2017), and Zhu et al. (2020). Condition (C2) requires the uniform boundedness of features $\mathbb{X}$, the weighting matrix $W$, and $S_{0}^{-1}$; this is a classical regularity condition in the SAR model (Lee, 2004; Yang and Lee, 2017). Condition (C3) is a typical assumption for the noise term (Huang et al., 2019; Zhu et al., 2020; Huang et al., 2020). It is remarkable that the artificially created noise is allowed to be non-Gaussian, as long as the moment conditions are satisfied. Furthermore, condition $E(e_{i}^{3})=E(\varepsilon_{i}^{3})=E(\varepsilon_{ij}^{x3})=0$ is assumed for convenience. The aim is to simplify the form of the asymptotic covariances for the estimators. Notably, the theory to be presented can be softly generalized under the assumption of non-zero third-order moments using the same theoretical proof techniques in this work. Given these conditions, we then establish the theoretical results in the subsequent subsections.
@@ 3.2. Theoretical Properties of the CLE
However, to estimate the parameters for the PSAR model based on the likelihood function, we need to analyze $\Omega^{-1}=\big{(}\sigma^{2}I_{N}+\lambda^{2}SS^{\top}\big{)}^{-1}$, which distinguishes the estimation procedure from the previous two approaches. On the one hand, even if $W$ and $S^{-1}$ satisfy the boundedness conditions of the $L_{1}$ and $L_{\infty}$ norms, it cannot be directly proven that $\Omega^{-1}$ is bounded in $L_{1}$ and $L_{\infty}$ norms. Consequently, the first method cannot be employed. On the other hand, notice that $S^{\top}S$ in $\Omega^{-1}$ cannot be analyzed as a transition-probability matrix; thus, the assumption of the existence of a stationary distribution cannot be made either. This makes the second method inapplicable. Therefore, we need to reexamine the theoretical properties of the CLE $\hat{\theta}_{\text{CLE}}$ based on the PSAR model. We begin with the discussion of the identification issue. Before presenting the theoretical results, the following identification condition is required.
@@
(Identification of CLE) Assume (C1)–(C4) hold. Then $\theta_{0}$ is identifiable.
The proof of Theorem 1 is given in Appendix C.1. We then examine the statistical property of the CLE, which results in the following theorem.
@@ 3.3. Discussion of Statistical Inference for the CLE
Remarkably, for the inference of $\hat{\theta}_{\text{CLE}}$, the asymptotic covariance in (3.1) cannot be immediately obtained when $\mathbb{X}$ is unobserved. Therefore, we rewrite $\Sigma_{k}^{\text{CL}}=\Sigma_{k}^{\text{CL}}(\mathbb{X},\theta)$ $(k=1,2)$ as a function of $\mathbb{X}$ and $\theta$. In this subsection, for simplicity, we use $\theta$ to represent the true parameter. Then, the difference between $\Sigma_{k}^{\text{CL}}(\mathbb{X}^{*},\theta)$ and $\Sigma_{k}^{\text{CL}}(\mathbb{X},\theta)$ should be considered. This is because data users can only observe $\mathbb{X}^{*}$ instead of $\mathbb{X}$, and statistical inference can be made relying only on $\mathbb{X}^{*}$. Thus, define $\Delta_{k}(\theta)=E\{\Sigma_{k}^{\text{CL}}(\mathbb{X}^{*},\theta)\}-\Sigma_{k}^{\text{CL}}(\mathbb{X},\theta)$. See Appendix A.1 for detailed expression of $\Delta_{k}(\theta)$. Therefore, based on the idea of bias correction, we can obtain a bias-corrected estimator for each element of the covariance matrix in (3.1) based on the observed $\mathbb{X}^{*}$. The estimators can be denoted as $\widehat{\Sigma}_{k}^{\text{CL}}=\Sigma_{k}^{\text{CL}}(\mathbb{X}^{*},\hat{\theta}_{\text{CLE}})-\Delta_{k}^{\text{CL}}(\hat{\theta}_{\text{CLE}})$ $(k=1,2)$ with $\hat{\theta}_{\text{CLE}}$ plugged in. Employing a technique similar to that used in the proof of Lemma 6 in Appendix B, the consistency of $\widehat{\Sigma}_{k}^{\text{CL}}$ can be established. Then, the estimated standard error of $\hat{\theta}_{\text{CLE}}$ could be calculated by the estimated asymptotic covariance matrix $(\widehat{\Sigma}_{2}^{\text{CL}})^{-1}\widehat{\Sigma}_{1}^{\text{CL}}(\widehat{\Sigma}_{2}^{\text{CL}})^{-1}$.
@@ 3.4. Theoretical Properties of the CLS
Here, the identification issue is discussed in two cases: with covariates and without covariates. We then have the following theorem.
@@
The proof of Theorem 4 is given in Appendix C.4. Next, we establish the following numerical and statistical theoretical properties of $\hat{\gamma}_{\text{CLS}}$.
@@
(Numerical Convergence and Asymptotic Normality of CLS) Assume that (C1)–(C3) and (C4) hold. If the initial value $\hat{\gamma}_{\rm CLS}^{(0)}$ lies close to $\hat{\gamma}_{\rm CLS}$, then $\hat{\gamma}_{\rm CLS}^{(t)}\to\hat{\gamma}_{\rm CLS}$ as $t\to\infty$ with probability tending to 1. And we further have $\sqrt{N}(\hat{\gamma}_{\rm CLS}-\gamma_{0})\to_{d}N(\mathbf{0}_{p+1},(\Sigma_{2}^{\rm LS})^{-1}\Sigma_{1}^{\rm LS}(\Sigma_{2}^{\rm LS})^{-1})$ as $N\to\infty$, where $\Sigma_{1}^{\rm LS}$ and $\Sigma_{2}^{\rm LS}$ are assumed to be positive definite matrices expressed as